### v1

# On the continuous properties for the 3D incompressible rotating Euler equations

Abstract: In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force.
Keywords: Euler equations; Coriolis force; Ill-posedness.
MSC (2010): 35B30; 76U05; 76B03.
1 Introduction
We say that the Cauchy problem
$\begin{cases}\partial_{t}f=F(f),\\ f(0,x)=f_{0}(x).\end{cases}$ (1.1)
is locally Hadamard well-posed in a Banach space $X$ if the following conditions hold
  1. 1.
    (Local existence)  For any initial data $u_{0}\in X$, there exists a short time $T=T(u_{0})>0$ and a solution $\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T),X)$ to the Cauchy problem (1.1);
  2. 2.
    (Uniqueness)  This solution $\mathbf{S}_{t}(u_{0})$ is unique in the space $\mathcal{C}([0,T),X)$;
  3. 3.
    (Continuous Dependence)  The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is continuous in the following sense: for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon,u_{0},\widetilde{u}_{0})>0$, such that if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
    $\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
Furthermore, we say that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniform continuous if for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ independent of $u_{0}$ or $\widetilde{u}_{0}$ such that, if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
$\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
We would like to emphasize that, in this paper, we say that the Cauchy problem (1.1) is well-posed in a Banach space $X$ if it enjoys existence, uniqueness and uniform continuous dependence on the initial data. In particular, we are interested in the lack of continuity with respect to time variable or uniform continuity of solution with respect to the initial data.
1.1 Known Well/Ill-posedness results
In this paper, we focus on the Cauchy problem for the Euler equations with the Coriolis force in $\mathbb{R}^{3}$, describing the motion of perfect incompressible fluids in the rotational framework:
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$ (1.2)
where the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible. The constant $\Omega\in\mathbb{R}$ represents the speed of rotation around the vertical unit vector $e_{3}=(0,0,1)$, which is called the Coriolis parameter. Problems concerning large-scale atmospheric and oceanic flows are known to be dominated by rotational effects. By this reason, almost all of the models of oceanography and meteorology dealing with large-scale phenomena include a Coriolis force. For example, an oceanic circulation featuring a hurricane is caused by the large rotation. There is no doubt that other physical effects are of similar significance like salinity, natural boundary conditions and so on. The dispersive effect of rotation in fluid flows has been studied in the literature from various perspectives, see, e.g., geophysical flows [16, 27, 31], life span and asymptotic behaviour in the case of fast rotations [10, 11, 15, 26] and almost global stability [17].
Let us first review some of the previous works in this direction.
The classical Euler equations.  In the case of $\Omega=0$, (1.2) reduces to the original Euler equations. Kato [21] proved the local well-posedness of classical solution to Euler equations in the Sobolev space $H^{s}(\mathbb{R}^{3})$ for all $s>5/2$. Kato-Ponce [23] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of the fractional order for $s>3/p+1,1<p<\infty$. Chae [7, 8, 9] and Chen-Miao-Zhang [12] gave further extensions to the Triebel-Lizorkin spaces $F_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1,1<p,q<\infty$ and the Besov spaces $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1<p<\infty,1\leq q\leq\infty$ or $s=3/p+1,1<p<\infty,q=1$. However, these two kinds of function spaces are only in the $L^{p}(1<p<\infty)$-framework since the Riesz transform is not bounded on $L^{\infty}$. The currently-known best result on the local existence was given by Pak-Park [30] in the Besov space $B_{\infty,1}^{1}(\mathbb{R}^{3})$. Guo-Li-Yin [18] proved the continuous dependence of the Euler equations in the space $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1\leq p\leq\infty,1\leq q<\infty$ or $s=3/p+1,1\leq p\leq\infty,q=1$. Later, Cheskidov-Shvydkoy [14] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{r,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<r\leq\infty$ and $s>d(2/r-1)$ if $1\leq r\leq 2$. Furthermore, Bourgain-Li [4, 5] proved the strong local ill-posedness of the Euler equations in borderline Besov spaces $B^{d/p+1}_{p,r}$ with $(p,r)\in[1,\infty)\times(1,\infty]$ when $d=2,3$. Subsequently, Misiołek-Yoneda [28] studied the borderline cases and showed that the 2D Euler equations are not locally well-posed in the sense of Hardamard in the $C^{1}$ space and in the Besov space $B^{1}_{\infty,1}$. Recently, Misiołek-Yoneda [29] showed that the solution map for the Euler equations is not even continuous in the space of Hölder continuous functions and thus not locally Hadamard well-posed in $B^{1+s}_{\infty,\infty}$ with any $s\in(0,1)$. Concerning the non-uniform continuity of the data-to-solution map, we would like to mention that the beautiful results of Himonas-Misiołek [19] covered both the torus $\mathbb{T}^{d}$ and the whole spaces $\mathbb{R}^{d}$ cases. More precisely, they proved that the solution map for the Euler equations in bi(tri)-dimension is not uniformly continuous in Sobolev spaces $H^{s}(\mathbb{T}^{d})$ for $s\in\mathbb{R}$ and in $H^{s}(\mathbb{R}^{d})$ for any $s>0$. Bourgain and Li [6] settled the border line case $s=0$.
The Euler equations with the Coriolis force.  In the case of $\Omega\neq 0$, as it is already observed in [10, 11, 15], the Euler-Coriolis system (1.2) exhibits a dispersion phenomenon which is due to the presence of the Coriolis force $\Omega e_{3}\times u$. For large Coriolis parameter $|\Omega|$, Dutrifoy [15] showed the asymptotics of solutions to vortex patches or Yudovich solutions as the Rossby number goes to zero for some particular initial data, and gave a lower bound on the lifespan $T_{\Omega}$ of the solution to (1.2) as $T_{\Omega}\gtrsim\log\log|\Omega|$. For $u_{0}\in H^{s}(\mathbb{R}^{3})$ with $s>5/2$, Koh-Lee-Takada [25] proved that there exists a unique local in time solution $u$ for (1.2) with $\Omega\in\mathbb{R}$ in the class $\mathcal{C}([0,T];H^{s}(\mathbb{R}^{3}))\cap\mathcal{C}^{1}([0,T];H^{s-1}(\mathbb{R}^{3}))$ (see also [33, 34]). Moreover, assuming that $s>7/2$, they showed that their solutions can be extended to long-time intervals $\left[0,T_{\Omega}\right]$ provided that the speed of rotation is large enough. V. Angulo-Castillo and L.C.F Ferreiragave [1] further gave extensions to the critical Besov space $B^{5/2}_{2,1}(\mathbb{R}^{3})$. Jia-Wan [20] proved the long time existence of classical solutions to (1.2) for initial data in the Sobolev space $H^{s}$ with $s>5/2$ by a weaker assumption on the lower bound of $|\Omega|$. However, the above mentioned results for (1.2) only involve the short-time existence, uniqueness for any speed of rotation and long-time existence for high speed of rotation. We should mention that the dispersion relation underlying the linear system
$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$
is given by $\Lambda(\xi)=\Omega\frac{\xi_{3}}{|\xi|}$. Compared with the classical Euler equations, we would like to emphasize that, this dispersive mechanism exhibits $O(t^{-1})$ decay rate in $L^{\infty}$-norm (see [32]) and is strongly anisotropic and degenerate. It is still surprising that, in the absence of viscosity, the Coriolis force alone is sufficient to stabilize the solutions globally in time in the full 3D setting (see [17]). It is found that the strong rotational effect enhances the temporal decay rate of a certain norm of the velocity (see [2]).
1.2 Motivation
A fundamental challenge in mathematical physics is to understand the behavior of solutions for the complex rotating fluids. From the PDE’s point of view, it is crucial to know if an equation which models a physical phenomenon is well-posed in the Hadamard’s sense: existence, uniqueness, and continuous dependence of the solutions with respect to the initial data. In particular, the lack of continuous dependence would cause incorrect solutions or non meaningful solutions. It is known that, the non-uniform continuity of data-to-solution map suggests that the local well-posedness cannot be established by the contraction mappings principle since this would imply Lipschitz continuity for the solution map. For certain types of evolutionary equations, for example, the Korteweg-de Vries (KdV) and nonlinear Schrödinger equations, the flow map is Lipschitz continuous. Paticularly, the KdV equation is shown to be well-posed in $H^{s}(\mathbb{R})$ with $s>-3/4$, then the solution map is Lipschitz on the same $H^{s}(\mathbb{R})$ (see [24]). However, for the Burgers equation $\partial_{t}u+u\partial_{x}u=0,$ which can be the most simple quasilinear symmetric hyperbolic equation, Kato [22] showed that the dependence on initial data of the Burgers equation in $H^{k}$ with $k\geq 2$ is continuous but not Hölder continuous with any prescribed exponent. This example indicates that the continuity of the solution map for a symmetric hyperbolic system is a delicate issue. Chen-Liu-Zhang [13] proved that the solution map of the b-family equation is Hölder continuous as a map from a bounded set of $H^{s}(\mathbb{R}),s>3/2$ with $H^{r}(\mathbb{R})(0\leq r<s)$ topology, to $C([0,T],H^{r}(\mathbb{R}))$ for some $T>0$. In this paper, for the Cauchy problem (1.2), we shall study the continuity of solution with respect to time variable or uniform continuity of solution with respect to the initial data.
Assume that $u_{0}\in B^{s}_{p,r}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Euler-Coriolis system (1.2). Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$. Therefore, we are interested in the following question:
$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
If the initial data $u_{0}$ have more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$. In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Euler-Coriolis system (1.2) can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.
1.3 Main Result
We can now state our main result as follows.
Theorem 1.1.
Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying
$s>\frac{3}{p}+1,(p,r)\in(1,\infty)\times[1,\infty]\quad\mathrm{or}\quad s=\frac{3}{p}+1,(p,r)\in(1,\infty)\times\{1\}.$ (1.3)
Then there exists a short time $T>0$ such that the Euler-Coriolis system (1.2) admits a unique solution $u\in L^{\infty}([0,T];B^{s}_{p,r})$ satisfying that
$\|u(t)\|_{L^{\infty}_{T}B^{s}_{p,r}}\lesssim\|u_{0}\|_{B^{s}_{p,r}}.$ (1.4)
Furthermore, if $r<\infty$, the above solution $u\in C([0,T];B^{s}_{p,r})$ is continuous dependence in $B^{s}_{p,r}$.
Remark 1.1.
Following the proof of the local well-posedness results for the Euler equations (see [3, Theorem 7.1 ] and [18, Theorem 1.1 ]), we can prove Theorem 1.1. Since the procedure is standard, we omit it.
Next, we consider the the property of continuous dependence of solutions of the Cauchy problem for (1.2) and show that the solution map for the Euler-Coriolis system (1.2) cannot be uniformly continuous in Besov spaces $B^{s}_{p,r}$. From now on, we denote
[…]
Theorem 1.2.
Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3). The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) is not uniformly continuous from any bounded subset in $B^{s}_{p,r}$ into $\mathcal{C}([0,T];B^{s}_{p,r})$. More precisely, there exist two initial data sequences $f_{n}+g_{n}$ and $f_{n}$ such that
$\left\|(f_{n})^{(2)}\right\|_{B^{s}_{p,r}}\approx 1\quad\text{and}\quad\lim_{n\rightarrow\infty}\left\|(f_{n})^{(1)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\left\|(f_{n})^{(3)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\|g_{n}\|_{B^{s}_{p,r}}=0$
but the distance of second component of the difference between the solution sequences $\mathbf{S}_{t}(f_{n}+g_{n})$ and $\mathbf{S}_{t}(f_{n})$
$\liminf_{n\rightarrow\infty}\left\|\left(\mathbf{S}_{t}(f_{n}+g_{n})-\mathbf{S}_{t}(f_{n})\right)^{(2)}\right\|_{B^{s}_{p,r}}\gtrsim t,\quad\forall\;t\in[0,T].$
Remark 1.2.
We should mention that Theorem 1.2 is stronger than that for the Euler equations in [19]. On the one hand, the non-uniform continuous is established in the more framework of Besov spaces. On the other hand, Theorem 1.2 is focused on the second component of data-to-solution map which leads to the non-uniform continuous.
Moreover, we obtain the following
Theorem 1.3.
Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3). For any $\alpha\in(0,1)$, there exits $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfying
$\limsup_{t\to 0^{+}}\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}=+\infty.$
Remark 1.3.
We would like to mention that Theorem 1.3 is new. In fact, Theorem 1.1 tells us that the solution $\mathbf{S}_{t}(u_{0})$ for (1.2) is continuous in time in Besov spaces $B^{s}_{p,r}$ with $r<\infty$ while Theorem 1.3 furthermore indicates that the solution $\mathbf{S}_{t}(u_{0})$ for (1.2) cannot be Hölder continuous in time in the same Besov spaces $B^{s}_{p,r}$.
Finally, as a by-product, we have
Theorem 1.4.
Assume that $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ with $(s,p,r=\infty)$ satisfying (1.3). There exits $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) satisfies
$\limsup_{t\to 0^{+}}\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,\infty}}\geq\varepsilon_{0}>0.$
Remark 1.4.
Theorem 1.4 implies the ill-posedness of (1.2) in the weaker Besov spaces $B^{s}_{p,\infty}(\mathbb{R}^{3})$ in the sense that the solution map to this system starting from $u_{0}$ is discontinuous at $t=0$ in the metric of $B^{s}_{p,\infty}$.
Remark 1.5.
As mentioned above, (1.2) with $\Omega=0$ is the classical Euler equations. Thus Our main Theorems 1.1-1.4 holds true for the classical Euler equations.
1.4 Strategies to the proof
We comment on a few points of main ideas and difficulties.
Proof of Theorem 1.1. In the case when $p\in(1,\infty)$, since the Coriolis force has no impact on the well-posedness results, we can prove Theorem 1.1 by following the proof of the local well-posedness results for the Euler equations (see [3, Theorem 7.1 ] and [18, Theorem 1.1 ]).
Proof of Theorem 1.2. The main task is to construct initial data such that the second component of the initial velocity is uniformly bounded in $B^{s}_{p,r}$, and in the same time to have the other component of the initial velocity remain small. We first choose two sequences of initial values $u^{n}_{0,1}=f_{n}+g_{n}$ and $u^{n}_{0,2}=f_{n}+g_{n}$, which can generate the approximate solutions $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$, respectively. Next, we perform perturbation analysis (See Proposition 3.1) to show the second component of the difference between the data-to-solution maps $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$ is bounded below by a positive constant at any later time, which means the solution map is not uniformly continuous.
Proof of Theorem 1.3. In order to prove the failure of Hölder regularity of solutions for the (1.2), we would like to first find initial condition to generate the corresponding solutions to (1.2) satisfying that some positive constant $C$
$\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}\geq Cn.$
Our key idea is to decompose the difference between the data-to-solution maps $\mathbf{S}_{t}(u_{0})$ and $\mathbf{S}_{t}(u_{0})$ as
$\mathbf{S}_{t}(u_{0})-u_{0}=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})-t\mathbf{v}_{0}(u_{0})\quad\text{and}\quad\mathbf{v}_{0}(u_{0})=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0}).$
There remains two main difficulties that we need to overcome.
  • First, we perform perturbation analysis to prove that $\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})$ is small;
  • Second, we need to ensure that the convect term $u_{0}\cdot\nabla u_{0}$ is large.
To carry out the analysis for these above two points, a big part of the technical difficulties lies in the construction of our initial data. We construct a new initial data which is somewhat different from Theorem 1.2. Finally, by utilizing the commutator estimate and some basic analysis, we hope to obtain the the loss of Hölder regularity of solutions to (1.2).
Proof of Theorem 1.4. The main point is to modify the initial data constructed in Theorem 1.3. Following the same procedure as that in Theorem 1.3, we can prove Theorem 1.4.
1.5 Organization of our paper
The paper is divided as follows.
  • In Section 2, we list some notations and known results which will be used in the sequel.
  • In Section 3, we show the non-uniform continuous of the solution map (Theorem 1.2).
  • In Section 4, we prove the failure of Hölder regularity of the solution map (Theorem 1.3).
  • In Section 5, we establish the discontinuous at zero time of the solution map (Theorem 1.4).
2 Preliminaries
We will use the following notations throughout this paper.
  • For $X$ a Banach space and $I\subset\mathbb{R}$, we denote by $\mathcal{C}(I;X)$ the set of continuous functions on $I$ with values in $X$. Sometimes we will denote $L^{p}(0,T;X)$ by $L_{T}^{p}X$.
  • The symbol $\mathrm{A}\lesssim(\gtrsim)\mathrm{B}$ means that there is a uniform positive “harmless” constant $\mathrm{C}$ independent of $\mathrm{A}$ and $\mathrm{B}$ such that $\mathrm{A}\leq(\geq)\mathrm{C}\mathrm{B}$, and we sometimes use the notation $\mathrm{A}\approx\mathrm{B}$ means that $\mathrm{A}\lesssim\mathrm{B}$ and $\mathrm{B}\lesssim\mathrm{A}$.
  • Let us recall that for all $u\in\mathcal{S}^{\prime}$, the Fourier transform $\mathcal{F}u$, also denoted by $\widehat{u}$, is defined by
    $\mathcal{F}u(\xi)=\widehat{u}(\xi)=\int_{\mathbb{R}^{3}}e^{-\mathrm{i}x\cdot\xi}u(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}^{3}.$
  • The inverse Fourier transform allows us to recover $u$ from $\widehat{u}$:
    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{3}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
  • We denote the projection
    $\mathcal{P}:L^{p}(\mathbb{R}^{3})\rightarrow L_{\sigma}^{p}(\mathbb{R}^{3})\equiv\overline{\left\{f\in\mathcal{C}^{\infty}_{0}(\mathbb{R}^{d});{\rm{div}}f=0\right\}}^{\|\cdot\|_{L^{p}(\mathbb{R}^{3})}},\quad p\in(1,\infty),$
    $\mathcal{Q}=\mathrm{Id}-\mathcal{P}.$
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant 3}$, where $\mathcal{P}_{ij}\equiv\delta_{ij}+R_{i}R_{j}$ with $\delta_{ij}$ being the Kronecker delta ($\delta_{ij}=0$ for $i\neq j$ and $\delta_{ii}=0$) and $R_{i}$ being the Riesz transform with symbol $-\mathrm{i}\xi_{1}/|\xi|$. Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
Next, we will recall some facts about the Littlewood-Paley decomposition and the nonhomogeneous Besov spaces (see [3] for more details). Choose a radial, non-negative, smooth function $\vartheta:\mathbb{R}^{3}\mapsto[0,1]$ such that ${\rm{supp}}\;\vartheta\subset B(0,4/3)$ and $\vartheta(\xi)\equiv 1$ for $|\xi|\leq 3/4$. Setting $\varphi(\xi):=\vartheta(\xi/2)-\vartheta(\xi)$, then we deduce that $\varphi$ has the following properties
  • ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}^{3}:3/4\leq|\xi|\leq 8/3\right\}$;
  • $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2;
  • $\vartheta(\xi)+\sum_{j\geq 0}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}^{3}$;
  • $\sum_{j\in\mathbb{Z}}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}^{3}\setminus\{0\}$.
The nonhomogeneous and homogeneous dyadic blocks are defined as follows
$\forall\,u\in\mathcal{S^{\prime}}(\mathbb{R}^{3}),\quad\Delta_{j}u=0,\;\text{if}\;j\leq-2;\quad\Delta_{-1}u=\vartheta(D)u;\quad\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\geq 0,$
and
$\forall\,u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{3}),\quad\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
where the pseudo-differential operator is defined by $\sigma(D):u\to\mathcal{F}^{-1}(\sigma\mathcal{F}u)$ and $\mathcal{S}^{\prime}_{h}$ is given by
$\mathcal{S}^{\prime}_{h}:=\Big{\{}u\in\mathcal{S^{\prime}}(\mathbb{R}^{3}):\;\lim_{j\rightarrow-\infty}\|\vartheta(2^{-j}D)u\|_{L^{\infty}}=0\Big{\}}.$
We recall the definition of the Besov Spaces and norms.
Definition 2.1 ([3]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the nonhomogeneous and homogeneous Besov spaces, respectively,
$B^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}:\;\|f\|_{B^{s}_{p,r}}:=\left\|2^{js}\|\Delta_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\geq-1)}<\infty\right\}$
and
$\dot{B}^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}_{h}:\;\|f\|_{\dot{B}^{s}_{p,r}}:=\left\|2^{js}\|\dot{\Delta}_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\in\mathbb{Z})}<\infty\right\}.$
We remark that, for any $s>0$ and $(p,r)\in[1,\infty]^{2}$, then $B^{s}_{p,r}(\mathbb{R}^{3})=\dot{B}^{s}_{p,r}(\mathbb{R}^{3})\cap L^{p}(\mathbb{R}^{3})$ and
$\|f\|_{B^{s}_{p,r}}\approx\|f\|_{L^{p}}+\|f\|_{\dot{B}^{s}_{p,r}}.$
The following Bernstein’s inequalities will be used in the sequel.
Lemma 2.1 ([3]).
Let $\mathcal{B}$ be a ball and $\mathcal{C}$ be an annulus. There exists a constant $C>0$ such that for all $k\in\mathbb{N}\cup\{0\}$, any $\lambda\in\mathbb{R}^{+}$ and any function $f\in L^{p}$ with $1\leq p\leq q\leq\infty$, we have
${\rm{supp}}\ \widehat{f}\subset\lambda\mathcal{B}\;\Rightarrow\;\|D^{k}f\|_{L^{q}}\leq C^{k+1}\lambda^{k+(\frac{1}{p}-\frac{1}{q})}\|f\|_{L^{p}},$
${\rm{supp}}\ \widehat{f}\subset\lambda\mathcal{C}\;\Rightarrow\;C^{-k-1}\lambda^{k}\|f\|_{L^{p}}\leq\|D^{k}f\|_{L^{p}}\leq C^{k+1}\lambda^{k}\|f\|_{L^{p}}.$
Next we recall the following product law which will be used often in the sequel.
Lemma 2.2 ([3]).
Assume $(s,p,r)$ satisfies (1.3) and $\sigma>0$. Then there exists a constant $C$, depending only on $d,p,r,\sigma$ or $s$ such that
$\|fg\|_{B^{\sigma}_{p,r}}\leq C\left(\|f\|_{L^{\infty}}\|g\|_{B^{\sigma}_{p,r}}+\|g\|_{L^{\infty}}\|f\|_{B^{\sigma}_{p,r}}\right),\quad\forall f,g\in L^{\infty}\cap B^{\sigma}_{p,r}.$
Furthermore, due to the embedding $B^{s-1}_{p,r}(\mathbb{R}^{3})\hookrightarrow L^{\infty}(\mathbb{R}^{3})$, there holds
$\|f\cdot\nabla g\|_{B^{s-1}_{p,r}}\leq C\|f\|_{B^{s-1}_{p,r}}\|g\|_{B^{s}_{p,r}},\quad\forall(f,g)\in B^{s-1}_{p,r}\times B^{s}_{p,r}$
and
$\|f\cdot\nabla g\|_{B^{s}_{p,r}}\leq C\left(\|f\|_{B^{s-1}_{p,r}}\|g\|_{B^{s+1}_{p,r}}+\|f\|_{B^{s}_{p,r}}\|g\|_{B^{s}_{p,r}}\right),\quad\forall(f,g)\in B^{s}_{p,r}\times B^{s+1}_{p,r}.$ (2.5)
Lemma 2.3 ([18]).
Assume that $(s,p,r)$ satisfies (1.3). Then there exists a constant $C$, depending only on $d,p,r,s$, such that for all $u,v\in B^{s}_{p,r}$ with $\mathrm{div\,}u=\mathrm{div\,}v=0$,
$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\|u\|_{B^{s}_{p,r}}\|v\|_{B^{s}_{p,r}}$
and
$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s-1}_{p,r}}\leq C\min\left\{\|u\|_{B^{s-1}_{p,r}}\|v\|_{B^{s}_{p,r}},\,\|v\|_{B^{s-1}_{p,r}}\|u\|_{B^{s}_{p,r}}\right\}.$
3 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Letting $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
${\widehat{\phi}(\xi)=}\begin{cases}1,&\text{if}\quad|\xi|\leq\frac{1}{64},\\ 0,&\text{if}\quad|\xi|\geq\frac{1}{8}.\end{cases}$
We assume that $\widehat{\phi}$ is even and real-valued such that $\theta$ is a real-valued function. Then we can define the new real-valued function
$\theta(x)=\phi(x_{1})\phi(x_{2})\phi(x_{3}).$
It is easy to check that $\theta(0)=\phi^{3}(0)>0$ and
$\mathrm{supp}\ \widehat{\theta}\subset\left\{\xi=(\xi_{1},\xi_{2},\xi_{3})\in\mathbb{R}^{3}:\ |\xi|\leq\frac{1}{2}\right\}.$ (3.6)
Thus for $\lambda\gg 1$ and $T\in\{\sin,\,\cos\}$
$\mathrm{supp}\ \mathcal{F}\Big{(}\theta(x)T(\lambda x_{1})\Big{)}\subset\left\{\xi\in\mathbb{R}^{3}:\ \lambda-\frac{1}{2}\leq|\xi|\leq\lambda+\frac{1}{2},\ |\xi_{2}|\leq\frac{1}{2},\ |\xi_{3}|\leq\frac{1}{2}\right\}.$ (3.7)
For any $p\in[1,\infty]$, there exist two positive constants $C_{1}$ and $C_{2}$ such that
$C_{1}\leq\|\phi\|_{L^{p}(\mathbb{R})}\leq C_{2}.$
Definition 1.  Let $(s,p,r)$ satisfies (1.3) and $n\gg 1$. We define the high frequency function $f_{n}$ and the low frequency function $g_{n}$ by
$f_{n}(x)=2^{-n(s+1)}\left(-\partial_{2},\partial_{1},0\right)\left[\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right)\right],$
$g_{n}(x)=2^{-n}\left(-\partial_{2},\partial_{1},0\right)\theta(x).$
Lemma 3.1.
Let $f_{n}$ and $g_{n}$ be defined as above. There exists a positive constant $C$ such that for any $\sigma\in\mathbb{R}$
$\|f_{n}^{(1)}\|_{B^{\sigma}_{p,r}}\leq C2^{n(\sigma-s-1)},$
$\|f_{n}^{(2)}\|_{B^{\sigma}_{p,r}}\thickapprox 2^{n(\sigma-s)},$
$\|f_{n}\|_{B_{p,r}^{\sigma}}\thickapprox 2^{n(\sigma-s)},$
$\|g_{n}\|_{B_{p,r}^{\sigma}}\thickapprox 2^{-n}.$
Proof.
From (3.7) and Bernstein’s inequality (see Lemma 2.1), it follows that the desired results of Lemma 3.1. ∎
Lemma 3.2.
Let $f_{n}$ and $g_{n}$ be defined as above. There exists a positive constant $c$ such that
$\liminf_{n\rightarrow\infty}\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,\infty}}\geq c.$ (3.8)
Proof.
Due to (3.6) and (3.7), we have
$\mathrm{supp}\ \mathcal{F}\left(g_{n}\cdot\nabla f_{n}^{(2)}\right)\subset\left\{\xi\in\mathbb{R}^{3}:\ \frac{17}{12}2^{n}-1\leq|\xi|\leq\frac{17}{12}2^{n}+1\right\},$
which implies
${\Delta_{j}\left(g_{n}\cdot\nabla f_{n}^{(2)}\right)=}\begin{cases}g_{n}\cdot\nabla f_{n}^{(2)},&\text{if}\;j=n,\\ 0,&\text{if}\;j\neq n.\end{cases}$
Thus, we have
$\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,\infty}}$ $=2^{ns}\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{L^{p}}.$ (3.9)
The definitions of $f_{n}$ and $g_{n}$ tell directly us that
$2^{ns}g_{n}\cdot\nabla f_{n}^{(2)}$ $=\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad+2^{-n}\phi(x_{1})\phi^{\prime}(x_{1})\sin\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad-2^{-2n}\phi(x_{1})\phi^{\prime\prime}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad+2^{-2n}\left[\phi^{\prime}(x_{1})\right]^{2}\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3}).$
Then, we have
$2^{ns}\|g_{n}\cdot\nabla f_{n}^{(2)}\|_{L^{p}}\geq\left\|\phi^{2}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(\mathbb{R})}\|\phi\|^{2}_{L^{2p}(\mathbb{R})}\left\|\phi(x)\phi^{\prime}(x)\right\|_{L^{p}(\mathbb{R})}-C2^{-n},$
Plugging the above into (3.9) yields
$\|g_{n}\cdot\nabla f_{n}\|_{B^{s}_{p,\infty}}$ $\geq c\left\|\phi^{2}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(\mathbb{R})}-C2^{-n}.$ (3.10)
Notice that the fact
$\liminf_{n\rightarrow\infty}\left\|\phi^{2}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(\mathbb{R})}\geq c>0,$
from (3.10), we get the desired result (3.8). Thus we finish the proof of Lemma 3.2.∎
Proposition 3.1.
Under the assumptions of Theorem 1.2, we have
$\|\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}\leq Ct^{2}\mathbf{E}(u_{0}),$ (3.11)
here and in what follows we denote $\mathbf{v}_{0}(u_{0}):=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0})$ and
$\mathbf{E}(u_{0})$ $:=1+\|u_{0}\|_{B^{s-1}_{p,r}}\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$
Proof.
For simplicity, we denote $u(t)=\mathbf{S}_{t}(u_{0})$ here and in what follows. By Theorem 1.1, one has
$\|u(t)\|_{B^{s+k}_{p,r}}\leq C\|u_{0}\|_{B^{s+k}_{p,r}}\quad\text{for }\;k\in\{0,\pm 1,2\}.$ (3.12)
Using the Mean Value Theorem and Lemma 2.2, we obtain from (1.2) that
$\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s-1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s-1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq C\int^{t}_{0}\left(\|u\|_{B^{s}_{p,r}}\|u\|_{B^{s-1}_{p,r}}+\|u\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\|u_{0}\|_{B^{s-1}_{p,r}},$ (3.13)
$\|u(t)-u_{0}\|_{B^{s}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\left(1+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+1}_{p,r}}\right),$ (3.14)
$\|u(t)-u_{0}\|_{B^{s+1}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s+1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s+1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s+1}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$ (3.15)
Using the Mean Value Theorem once again, we obtain that
$\|u(t)-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u+\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\|\mathcal{Q}(u)-\mathcal{Q}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\cdot\nabla u-u_{0}\cdot\nabla u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\quad+\Omega\int^{t}_{0}\|\mathcal{P}\left(e_{3}\times(u-u_{0})\right)\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\|u(\tau)\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
$\quad\ +\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\|u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\|u_{0}\|_{B^{s+1}_{p,r}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$\quad\ +\|u_{0}\|_{L^{\infty}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\mathrm{d}\tau.$ (3.16)
Plugging (3)-(3) into (3) and using the fact that $B^{s-1}_{p,r}(\mathbb{R}^{3})\hookrightarrow L^{\infty}(\mathbb{R}^{3})$, yields the desired result of Proposition 3.1. ∎
Proof of Theorem 1.2. Let $u^{n}_{0}:=f_{n}+g_{n}$. Obviously, we have
$\|u^{n}_{0}-f_{n}\|_{B^{s}_{p,r}}=\|g_{n}\|_{B^{s}_{p,r}}\leq C2^{-n},$
which means that
$\lim_{n\to\infty}\|u^{n}_{0}-f_{n}\|_{B^{s}_{p,r}}=0.$
Obviously, we obtain from Lemma 3.1 that
$\|u^{n}_{0}\|_{B^{s+k}_{p,r}}\leq C2^{kn}\quad\text{for }\;k\in\{0,\pm 1,2\}.$
We decompose the solution maps $\mathbf{S}_{t}(u^{n}_{0})$ and $\mathbf{S}_{t}(f_{n})$ as follows
$\mathbf{S}_{t}(u^{n}_{0})$ $=\mathbf{S}_{t}(u^{n}_{0})-u^{n}_{0}+t\mathbf{v}_{0}(u^{n}_{0})-(f_{n}+g_{n})-t(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})$
$\quad+t\mathcal{Q}\big{[}(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})\big{]}-t\mathcal{P}(\Omega e_{3}\times(f_{n}+g_{n})),$
$\mathbf{S}_{t}(f_{n})$ $=\mathbf{S}_{t}(f_{n})-f_{n}+t\mathbf{v}_{0}(f_{n})-f_{n}-tf_{n}\cdot\nabla f_{n}$
$\quad+t\mathcal{Q}(f_{n}\cdot\nabla f_{n})-t\mathcal{P}(\Omega e_{3}\times f_{n}).$
Then, by Proposition 3.1, we deduce that
$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$ $\geq t\left\|\left[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\right]^{(2)}\right\|_{B^{s}_{p,r}}-t\|\mathcal{P}(\Omega e_{3}\times g_{n})\|_{B^{s}_{p,r}}$
$\quad-t\left\|\mathcal{Q}\big{[}(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\big{]}\right\|_{B^{s}_{p,r}}-\|g_{n}\|_{B^{s}_{p,r}}-Ct^{2}$
$\geq t\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,r}}-Ct^{2}-C2^{-n},$ (3.17)
where we have used the facts from Lemma 2.3 and Lemma 3.1
$\|\mathcal{Q}(f_{n}\cdot\nabla g_{n})\|_{B^{s}_{p,r}}\lesssim\|f_{n}\cdot\nabla g_{n}\|_{B^{s}_{p,r}}\lesssim\|f_{n}\|_{B^{s}_{p,r}}\|g_{n}\|_{B^{s+1}_{p,r}}\lesssim 2^{-n},$
$\|\mathcal{Q}(g_{n}\cdot\nabla g_{n})\|_{B^{s}_{p,r}}\lesssim\|g_{n}\cdot\nabla g_{n}\|_{B^{s}_{p,r}}\lesssim\|g_{n}\|_{B^{s}_{p,r}}\|g_{n}\|_{B^{s+1}_{p,r}}\lesssim 2^{-2n},$
$\|\mathcal{Q}(g_{n}\cdot\nabla f_{n})\|_{B^{s}_{p,r}}=\|\mathcal{Q}(f_{n}\cdot\nabla g_{n})\|_{B^{s}_{p,r}}\lesssim\|f_{n}\|_{B^{s}_{p,r}}\|g_{n}\|_{B^{s+1}_{p,r}}\lesssim 2^{-n}.$
Using Lemma 3.2, we obtain from (3) that
$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$ $\geq ct-Ct^{2}-C2^{-n},$
which yields the result of Theorem 1.2. Thus we complete the proof of Theorem 1.2.
4 Proof of Theorem 1.3
Assume that $(s,p,r)$ satisfies (1.3). We define the initial data $u_{0}(x)$ as
$u_{0}(x):=\left(-\partial_{2},\partial_{1},0\right)\sum\limits^{\infty}_{n=3}a_{n}(x),\quad\text{where}$
$a_{n}(x):=n^{-2}2^{-n(s+1)}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
Lemma 4.1.
Assume that $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. Define the divergence-free vector field $u_{0}$ as above. Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some positive constant $c_{0}$ such that
$\|u_{0}\|_{B^{s}_{p,r}}\thickapprox\|u_{0}\|_{\dot{B}^{s}_{p,r}}\thickapprox 1,$ (4.18)
$2^{ns}\left\|u_{0}\cdot\nabla\Delta_{n}u_{0}^{(2)}\right\|_{L^{p}}\geq c_{0}n^{-2}2^{n}.$ (4.19)
Proof.
Due to the simple fact (3.7), it holds that for some sufficiently large $n\in\mathbb{Z}^{+}$
$\Delta_{n}u_{0}^{(1)}=-\partial_{2}a_{n}\quad\text{and}\quad\Delta_{n}u_{0}^{(2)}=\partial_{1}a_{n},$
which gives directly that
$u_{0}\cdot\nabla\Delta_{n}u_{0}^{(2)}$ $=-u_{0}^{(1)}{\partial_{1}\partial_{1}}a_{n}+u_{0}^{(2)}\partial_{1}\partial_{2}a_{n}.$ (4.20)
(4.18) is obvious. To estimate (4.19), we should be emphasize that the leading term is $u_{0}^{(1)}{\partial_{1}\partial_{1}}a_{n}$ since the other terms in the right hand side of (4.20) can be absorbed by the term $u_{0}^{(1)}{\partial_{1}\partial_{1}}a_{n}$. Thus we focus on the estimation of the leading term.
$-n^{2}2^{n(s+1)}u_{0}^{(1)}{\partial_{1}\partial_{1}}a_{n}$ $=\left(\frac{17}{12}2^{n}\right)^{2}\sum_{k=3}^{\infty}k^{-2}2^{-k(s+1)}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{k}x_{1}\right)$
×$\quad\times\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi^{2}(x_{3})\phi(x_{2})\phi^{\prime}(x_{2})+\text{Remainder terms}.$
For the Remainder terms, noticing that at most one partial derivative falls on cosine function, thus we find easily that
$\left\|\text{Remainder terms}\right\|_{L^{p}(\mathbb{R}^{3})}\leq C2^{n}.$
Since $\sum_{k=3}^{\infty}k^{-2}2^{-k(s+1)}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{k}x_{1}\right)$ is a real-valued and continuous function on $\mathbb{R}$, then there exists some $\delta>0$ such that for any $x\in B_{\delta}(0):=\{x\in\mathbb{R}:\;|x|\leq\delta\}$
$\left|\sum_{k=3}^{\infty}k^{-2}2^{-k(s+1)}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{k}x_{1}\right)\right|\geq\frac{\phi^{2}(0)}{2}\sum\limits^{\infty}_{k=3}k^{-2}2^{-k(s+1)}=:c_{0}>0.$
Thus we have for some sufficiently large $n\in\mathbb{Z}^{+}$
$n^{2}2^{n(s+1)}\left\|u_{0}^{(1)}{\bf\partial_{1}\partial_{1}}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\geq c_{0}2^{2n}-C2^{n}.$ (4.21)
By easy computations, we have
$n^{2}2^{n(s+1)}\left\|u_{0}^{(2)}\partial_{1}\partial_{2}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\leq C2^{n}.$ (4.22)
Noticing that (4.20), from (4.21) and (4.22), we obtain that for some sufficiently large $n\in\mathbb{Z}^{+}$
$n^{2}2^{n(s+1)}\left\|u_{0}\cdot\nabla\Delta_{k}u_{0}^{(2)}\right\|_{L^{p}(\mathbb{R}^{3})}\geq c_{0}2^{2n}-C2^{n},$
which is nothing but the desired result (4.19). We complete the proof of Lemma 4.1. ∎
Proposition 4.1.
Assume that $\|u_{0}\|_{B^{s}_{p,r}}\approx 1$. Under the assumptions of Theorem 1.3, we have
$\|\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{{B}^{s-2}_{p,r}}\lesssim t^{2}.$
Proof.
It holds that
$\|u(t)\|_{L^{\infty}_{T}(B^{s}_{p,r})}\lesssim\|u_{0}\|_{B^{s}_{p,r}}\lesssim 1.$
From (3), we obtain
$\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}\lesssim t\|u_{0}\|_{B^{s}_{p,r}}\lesssim t.$ (4.23)
By the Newton-Leibniz formula again, we get
$\|\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{{B}^{s-2}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u+\mathbf{v}_{0}\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\mathcal{P}\left(u\cdot\nabla u-u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times(u-u_{0})\right)\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\left(\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$
We complete the proof of Proposition 4.1. ∎
Proof of Theorem 1.3. Setting $\mathbf{w}:=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})$
$\mathbf{S}_{t}(u_{0})-u_{0}=-t\mathbf{v}_{0}(u_{0})+\mathbf{w}\quad\text{and}\quad\mathbf{v}_{0}(u_{0})=u_{0}\cdot\nabla u_{0}-\mathcal{Q}(u_{0}\cdot\nabla u_{0})+\Omega\mathcal{P}(e_{3}\times u_{0}).$
By the triangle inequality, we deduce that
$\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\left\|\Delta_{n}\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{L^{p}}=2^{{ns}}\left\|\Delta_{n}\big{(}t\mathbf{v}_{0}(u_{0})-\mathbf{w}\big{)}^{(2)}\right\|_{L^{p}}$
$\geq t2^{{ns}}\left\|\Delta_{n}\left(\mathbf{v}_{0}(u_{0})\right)^{(2)}\right\|_{L^{p}}-2^{{2n}}2^{{n(s-2)}}\left\|\Delta_{n}\mathbf{w}\right\|_{L^{p}}$
$\geq t2^{{n}s}\left\|\Delta_{n}\big{(}u_{0}\cdot\nabla u_{0}^{(2)}\big{)}\right\|_{L^{p}}-t2^{{n}s}\left\|\Delta_{n}\mathcal{Q}(u_{0}\cdot\nabla u_{0})\right\|_{L^{p}}$
$\quad-t2^{{n}s}\left\|\Delta_{n}\big{(}\Omega\mathcal{P}(e_{3}\times u_{0}))\big{)}\right\|_{L^{p}}-C2^{2{n}}\|\mathbf{w}\|_{\dot{B}^{s-2}_{p,\infty}}$
$\geq t2^{{n}s}\left\|u_{0}\cdot\nabla\Delta_{n}u_{0}^{(2)}\right\|_{L^{p}}-t2^{ns}\left\|[\Delta_{n},u_{0}]\cdot\nabla u_{0}^{(2)}\right\|_{L^{p}}$
$\quad-Ct2^{{n}s}\left\|\Delta_{n}\mathcal{Q}(u_{0}\cdot\nabla u_{0})\right\|_{L^{p}}-Ct\|u_{0}\|_{B^{s}_{p,\infty}}-C2^{2n}t^{2}$
$\geq t2^{ns}\left\|u_{0}\cdot\nabla\Delta_{n}u_{0}^{(2)}\right\|_{L^{p}}-Ct-C2^{2{n}}t^{2},$ (4.24)
where we have used the fact:
$\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\cdot\nabla u_{0}^{(2)}\|_{L^{p}}\big{\|}_{\ell^{\infty}}\lesssim\|\nabla u_{0}\|_{L^{\infty}}\|u_{0}\|_{B^{s}_{p,r}}\lesssim 1,$
and
$2^{{ns}}\|\Delta_{n}\mathcal{Q}(u_{0}\cdot\nabla u_{0})\|_{L^{p}}$ $=2^{{ns}}\|\Delta_{n}((-\Delta)^{-1}\nabla(\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}))\|_{L^{p}}$
$\approx 2^{{n(s-1)}}\|\Delta_{n}(\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i})\|_{L^{p}}$
$\lesssim\|\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}\|_{B^{s-1}_{p,\infty}}$
$\lesssim\|u_{0}\|^{2}_{B^{s}_{p,\infty}}\lesssim 1.$
Then, using Lemma 4.1, we obtain from (4) that
$\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,r}}\geq ctn^{-2}2^{n}-Ct-C2^{2{n}}t^{2},$
which implies
$t^{-\alpha}\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{n}-Ct^{1-\alpha}-C2^{2{n}}t^{2-\alpha},$
Thus, picking $t^{1-\alpha}_{n}=n^{3}2^{-n}$ with large $n$, we have
$t^{-\alpha}_{n}\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,r}}$ $\geq cn-Cn^{3}2^{-n}-Cn^{6}t_{n}^{\alpha}\geq\tilde{c}n.$
This completes the proof of Theorem 1.3.
5 Proof of Theorem 1.4
We define the initial data $u_{0}(x)$ as
$u_{0}(x):=\left(-\partial_{2},\partial_{1},0\right)\sum\limits^{\infty}_{n=3}b_{n}(x),\quad\text{where}$
$b_{n}(x):=2^{-ns}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
Following similar procedure as that in Theorem 1.3. Then, we have
$\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,\infty}}$ $\geq 2^{{ns}}\left\|\Delta_{n}\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{L^{p}}=2^{{ns}}\left\|\Delta_{n}\big{(}t\mathbf{v}_{0}(u_{0})-\mathbf{w}\big{)}^{(2)}\right\|_{L^{p}}$
$\geq t2^{{n}s}\left\|\Delta_{n}\big{(}u_{0}\cdot\nabla u_{0}^{(2)}\big{)}\right\|_{L^{p}}-t2^{{n}s}\left\|\Delta_{n}\mathcal{Q}(u_{0}\cdot\nabla u_{0})\right\|_{L^{p}}$
$\quad-t2^{{n}s}\left\|\Delta_{n}\big{(}\Omega\mathcal{P}(e_{3}\times u_{0}))\big{)}\right\|_{L^{p}}-C2^{2{n}}\|\mathbf{w}\|_{\dot{B}^{s-2}_{p,\infty}}$
$\geq ct2^{n}-Ct-C2^{2n}t^{2}.$
Thus, picking $t2^{kn}\approx\varepsilon$ with small $\varepsilon$, we have
$\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,\infty}}\geq c\varepsilon-C\varepsilon^{2}\geq\tilde{c}\varepsilon.$
This completes the proof of Theorem 1.4.
6 Discussion and Conclusion
In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force. In this paper, our contributions on the continuous properties for the 3D incompressible rotating Euler equationsare lies in three aspects: by constructing some proper, new and different initial conditions, we showed the non-uniform continuous of the solution map (Theorem 1.2), proved the failure of Hölder regularity of the solution map (Theorem 1.3) and established the discontinuous at $t=0$ of the solution map (Theorem 1.4).
Future directions. In the case when $p\in(1,\infty)$, since the Coriolis force term has no impact on the well-posedness results, we can prove Theorem 1.1 by following the proof of the local well-posedness results for the Euler equations (see [3, Theorem 7.1 ] and [18, Theorem 1.1 ]). However, in the case when $p=1$ or $p=\infty$, due to the the appearance of the dispersive effect of rotation which leads to the fact that Riesz transform does not map continuously from $L^{\infty}(\text{or}\;L^{1})$ to $L^{\infty}(\text{or}\;L^{1})$, we will encounter the main difficulty when establishing the uniform bounds of solution in $L^{\infty}(\text{or}\;L^{1})$-based spaces. We believe that, the dispersive effect of rotation is able to prevent well-posedness for the Cauchy problem (1.2). More precisely, we expect that the Cauchy problem (1.2) is ill-posed in $L^{\infty}(\text{or}\;L^{1})$-based Besov spaces. We will investigate this interesting problem in the future work.
Acknowledgments
J. Li is partially supported by the National Natural Science Foundation of China (No. 12161004), Innovative High end Talent Project in Ganpo Talent Program (No. gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20232BCJ23009). Y. Yu is partially supported by the National Natural Science Foundation of China (No. 12101011). N. Zhu is partially supported by the National Natural Science Foundation of China (No. 12161055), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20243BCE51079).
Declarations
Data Availability No data was used for the research described in the article.
Conflict of interest The authors declare that they have no conflict of interest.

### v2

@@
 Math.
-Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces.
+Phys., 296, 2010] to the more general framework of Besov spaces; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces.
 In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero.
@@
 1 Introduction
-We say that the Cauchy problem
-$\begin{cases}\partial_{t}f=F(f),\\ f(0,x)=f_{0}(x).\end{cases}$ (1.1)
+In this paper, we focus on the Cauchy problem for the Euler equations with the Coriolis force in $\mathbb{R}^{3}$, describing the motion of perfect incompressible fluids in the rotational framework:
+$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3},\end{cases}$ (1.1)
+where the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible.
+The constant $\Omega\in\mathbb{R}$ represents the speed of rotation around the vertical unit vector $e_{3}=(0,0,1)$ and is called the Coriolis parameter.
+Problems concerning large-scale atmospheric and oceanic flows are known to be dominated by rotational effects.
+For this reason, almost all of the models of oceanography and meteorology dealing with large-scale phenomena include a Coriolis force.
+For example, an oceanic circulation featuring a hurricane is caused by the large rotation.
+There is no doubt that other physical effects are of similar significance like salinity, natural boundary conditions and so on.
+The dispersive effect of rotation in fluid flows has been studied in the literature from various perspectives, see, e.g., geophysical flows [16, 27, 32], life span and asymptotic behaviour in the case of fast rotations [10, 11, 15, 26] and almost global stability [17].
+Before stating related results for (1.1) precisely, we recall the notion of well-posedness in the sense of Hadamard. We say that the Cauchy problem
+$\begin{cases}\partial_{t}f=F(f),\\ f(0,x)=u_{0}(x),\end{cases}$ (1.2)
 is locally Hadamard well-posed in a Banach space $X$ if the following conditions hold
@@
   3. 3.
-    (Continuous Dependence)  The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is continuous in the following sense: for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon,u_{0},\widetilde{u}_{0})>0$, such that if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
-    $\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
-Furthermore, we say that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniform continuous if for any $T_{1}<T$ and $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ independent of $u_{0}$ or $\widetilde{u}_{0}$ such that, if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
-$\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T),X)}\leq\varepsilon.$
-We would like to emphasize that, in this paper, we say that the Cauchy problem (1.1) is well-posed in a Banach space $X$ if it enjoys existence, uniqueness and uniform continuous dependence on the initial data.
+    (Continuous Dependence)  The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is continuous in the following sense: for any $u_{0}$, any $T_{1}<T\left(u_{0}\right)$ and any $\varepsilon>0$, there exists $\delta=\delta(\varepsilon,u_{0})$ such that for any $\widetilde{u}_{0}$, if $\left\|u_{0}-\widetilde{u}_{0}\right\|_{X}\leq\delta$ then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
+    $\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T_{1}],X)}\leq\varepsilon.$
+Furthermore, we say that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniform continuous if for any $T_{1}<T$ and any $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ independent of $u_{0}$ or $\widetilde{u}_{0}$ such that, if $\|u_{0}-\widetilde{u}_{0}\|_{X}\leq\delta$, then $\mathbf{S}_{t}(\widetilde{u}_{0})$ exists up to $T_{1}$ and
+$\|\mathbf{S}_{t}(u_{0})-\mathbf{S}_{t}(\widetilde{u}_{0})\|_{\mathcal{C}([0,T_{1}],X)}\leq\varepsilon.$
+We would like to emphasize that, in this paper, we say the Cauchy problem (1.2) is well-posed in a Banach space $X$ if it enjoys existence, uniqueness and continuous dependence on the initial data.
+We say that the Cauchy problem (1.2) is semi-linearly well-posed (see [30]) for the initial data in $X$ if it is Hadamard well-posed and in addition, the solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ is uniformly continuous from any bounded subset $B\subset X$ into $C\left([0,T];X\right)$.
 In particular, we are interested in the lack of continuity with respect to time variable or uniform continuity of solution with respect to the initial data.
 1.1 Known Well/Ill-posedness results
-In this paper, we focus on the Cauchy problem for the Euler equations with the Coriolis force in $\mathbb{R}^{3}$, describing the motion of perfect incompressible fluids in the rotational framework:
-$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+u\cdot\nabla u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$ (1.2)
-where the vector field $u(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ stands for the velocity of the fluid, the quantity $P(t,x):[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}$ denotes the scalar pressure, and $\mathrm{div\,}u=0$ means that the fluid is incompressible.
-The constant $\Omega\in\mathbb{R}$ represents the speed of rotation around the vertical unit vector $e_{3}=(0,0,1)$, which is called the Coriolis parameter.
-Problems concerning large-scale atmospheric and oceanic flows are known to be dominated by rotational effects.
-By this reason, almost all of the models of oceanography and meteorology dealing with large-scale phenomena include a Coriolis force.
-For example, an oceanic circulation featuring a hurricane is caused by the large rotation.
-There is no doubt that other physical effects are of similar significance like salinity, natural boundary conditions and so on.
-The dispersive effect of rotation in fluid flows has been studied in the literature from various perspectives, see, e.g., geophysical flows [16, 27, 31], life span and asymptotic behaviour in the case of fast rotations [10, 11, 15, 26] and almost global stability [17].
 Let us first review some of the previous works in this direction.
@@
 Kato [21] proved the local well-posedness of classical solution to Euler equations in the Sobolev space $H^{s}(\mathbb{R}^{3})$ for all $s>5/2$.
-Kato-Ponce [23] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of the fractional order for $s>3/p+1,1<p<\infty$.
+Kato-Ponce [23] extended this result to the Sobolev spaces $W^{s,p}(\mathbb{R}^{3})$ of fractional order for $s>3/p+1,1<p<\infty$.
 Chae [7, 8, 9] and Chen-Miao-Zhang [12] gave further extensions to the Triebel-Lizorkin spaces $F_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1,1<p,q<\infty$ and the Besov spaces $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1<p<\infty,1\leq q\leq\infty$ or $s=3/p+1,1<p<\infty,q=1$.
@@
 Guo-Li-Yin [18] proved the continuous dependence of the Euler equations in the space $B_{p,q}^{s}(\mathbb{R}^{3})$ with $s>3/p+1$, $1\leq p\leq\infty,1\leq q<\infty$ or $s=3/p+1,1\leq p\leq\infty,q=1$.
-Later, Cheskidov-Shvydkoy [14] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{r,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<r\leq\infty$ and $s>d(2/r-1)$ if $1\leq r\leq 2$.
+Later, Cheskidov-Shvydkoy [14] proved that the solution of the Euler equations cannot be continuous as a function of the time variable at $t=0$ in the spaces $B^{s}_{r,\infty}(\mathbb{T}^{d})$ where $s>0$ if $2<r\leq\infty$ and s>d(2/r1)s>d(2/r-1) if $1\leq r\leq 2$.
 Furthermore, Bourgain-Li [4, 5] proved the strong local ill-posedness of the Euler equations in borderline Besov spaces $B^{d/p+1}_{p,r}$ with $(p,r)\in[1,\infty)\times(1,\infty]$ when $d=2,3$.
@@
 V.
-Angulo-Castillo and L.C.F Ferreiragave [1] further gave extensions to the critical Besov space $B^{5/2}_{2,1}(\mathbb{R}^{3})$.
-Jia-Wan [20] proved the long time existence of classical solutions to (1.2) for initial data in the Sobolev space $H^{s}$ with $s>5/2$ by a weaker assumption on the lower bound of $|\Omega|$.
+Angulo-Castillo and L.C.F Ferreira [1] further gave extensions to the critical Besov space $B^{5/2}_{2,1}(\mathbb{R}^{3})$.
+Jia-Wan [20] proved the long time existence of classical solutions to (1.1) for initial data in the Sobolev space $H^{s}(s>5/2)$ with a weaker assumption on the lower bound of $|\Omega|$.
 However, the above mentioned results for (1.1) only involve the short-time existence, uniqueness for any speed of rotation and long-time existence for high speed of rotation.
 We should mention that the dispersion relation underlying the linear system
-$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3}.\end{cases}$
+$\begin{cases}\partial_{t}u+\Omega e_{3}\times u+\nabla P=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ \mathrm{div\,}u=0,&\quad(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3},\\ u(t=0,x)=u_{0}(x),&\quad x\in\mathbb{R}^{3},\end{cases}$
 is given by $\Lambda(\xi)=\Omega\frac{\xi_{3}}{|\xi|}$.
-Compared with the classical Euler equations, we would like to emphasize that, this dispersive mechanism exhibits $O(t^{-1})$ decay rate in $L^{\infty}$-norm (see [32]) and is strongly anisotropic and degenerate.
+Compared with the classical Euler equations, we would like to emphasize that this dispersive mechanism exhibits $O(t^{-1})$ decay rate in $L^{\infty}$-norm (see [33]) and is strongly anisotropic and degenerate.
 It is still surprising that, in the absence of viscosity, the Coriolis force alone is sufficient to stabilize the solutions globally in time in the full 3D setting (see [17]).
@@
 For certain types of evolutionary equations, for example, the Korteweg-de Vries (KdV) and nonlinear Schrödinger equations, the flow map is Lipschitz continuous.
-Paticularly, the KdV equation is shown to be well-posed in $H^{s}(\mathbb{R})$ with $s>-3/4$, then the solution map is Lipschitz on the same $H^{s}(\mathbb{R})$ (see [24]).
-However, for the Burgers equation $\partial_{t}u+u\partial_{x}u=0,$ which can be the most simple quasilinear symmetric hyperbolic equation, Kato [22] showed that the dependence on initial data of the Burgers equation in $H^{k}$ with $k\geq 2$ is continuous but not Hölder continuous with any prescribed exponent.
+Paticularly, the KdV equation is shown to be well-posed in $H^{s}(\mathbb{R})$ with s>3/4s>-3/4, then the solution map is Lipschitz on the same $H^{s}(\mathbb{R})$ (see [24]).
+However, for the Burgers equation $\partial_{t}u+u\partial_{x}u=0$, which is the most simple quasilinear symmetric hyperbolic equation, Kato [22] showed that the dependence on initial data of the Burgers equation in $H^{k}$ with $k\geq 2$ is continuous but not Hölder continuous with any prescribed exponent.
 This example indicates that the continuity of the solution map for a symmetric hyperbolic system is a delicate issue.
@@
 In this paper, for the Cauchy problem (1.1), we shall study the continuity of solution with respect to time variable or uniform continuity of solution with respect to the initial data.
-Assume that $u_{0}\in B^{s}_{p,r}$, it is known that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Euler-Coriolis system (1.2).
+Assume that $u_{0}\in B^{s}_{p,r}$, based on the local well-posedness theory for the Euler equations (see [3, Theorem 7.1]), we know that there exists a solution $u\in\mathcal{C}([0,T];B^{s}_{p,r})$ for the Euler-Coriolis system (1.1).
 Naturally, we may wonder whether or not the solution $u$ can belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with some $\alpha\in(0,1)$.
 Therefore, we are interested in the following question:
-$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists 1\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
-If the initial data $u_{0}$ have more regularity such that $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$.
+$u_{0}\in B^{s}_{p,r}\;\Rightarrow\;\exists\;u\in\mathcal{C}([0,T];B^{s}_{p,r})\;\overset{?}{\Rightarrow}\;u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})\quad\text{with}\;\alpha\in(0,1).$
+If the initial data $u_{0}$ has more regularity, that is $u_{0}\in B^{s^{\prime}}_{p,r}$ for some $s^{\prime}>s$, by the interpolation argument, we can deduce that $u\in\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with $\alpha=s^{\prime}-s$.
 In this paper, we will show that there exits initial data $u_{0}\in B^{s}_{p,r}$ such that the corresponding solution of the Euler-Coriolis system (1.1) can not belong to $\mathcal{C}^{\alpha}([0,T];B^{s}_{p,r})$ with any $\alpha\in(0,1)$.
@@
 $\|u(t)\|_{L^{\infty}_{T}B^{s}_{p,r}}\lesssim\|u_{0}\|_{B^{s}_{p,r}}.$ (1.4)
-Furthermore, if $r<\infty$, the above solution $u\in C([0,T];B^{s}_{p,r})$ is continuous dependence in $B^{s}_{p,r}$.
+Furthermore, if $r<\infty$, the above solution $u\in C([0,T];B^{s}_{p,r})$ is continuous dependent in $B^{s}_{p,r}$.
 Remark 1.1.
 Following the proof of the local well-posedness results for the Euler equations (see [3, Theorem 7.1] and [18, Theorem 1.1]), we can prove Theorem 1.1. Since the procedure is standard, we omit it.
-Next, we consider the the property of continuous dependence of solutions of the Cauchy problem for (1.2) and show that the solution map for the Euler-Coriolis system (1.2) cannot be uniformly continuous in Besov spaces $B^{s}_{p,r}$.
+Next, we consider the property of continuous dependence of solutions of the Cauchy problem for (1.1) and show that the solution map for the Euler-Coriolis system (1.1) cannot be uniformly continuous in Besov spaces $B^{s}_{p,r}$.
 From now on, we denote
 […]
+and denote any bounded subset $U_{R}$ in $B^{s}_{p,r}(\mathbb{R}^{3})$ by
+$U_{R}:=\left\{u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3}):\|u_{0}\|_{B^{s}_{p,r}(\mathbb{R}^{3})}\leq R,\;\mathrm{div}u_{0}=0\right\}.$
 Theorem 1.2.
-Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3).
-The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) is not uniformly continuous from any bounded subset in $B^{s}_{p,r}$ into $\mathcal{C}([0,T];B^{s}_{p,r})$.
+Assume that $(s,p,r)$ satisfying (1.3).
+The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.1) is not uniformly continuous from any bounded subset in $U_{R}\subset B^{s}_{p,r}$ into $\mathcal{C}([0,T];B^{s}_{p,r})$.
 More precisely, there exist two initial data sequences $f_{n}+g_{n}$ and $f_{n}$ such that
-$\left\|(f_{n})^{(2)}\right\|_{B^{s}_{p,r}}\approx 1\quad\text{and}\quad\lim_{n\rightarrow\infty}\left\|(f_{n})^{(1)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\left\|(f_{n})^{(3)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\|g_{n}\|_{B^{s}_{p,r}}=0$
+$\left\|(f_{n})^{(2)}\right\|_{B^{s}_{p,r}}\approx 1\quad\text{and}$
+$\lim_{n\rightarrow\infty}\left\|(f_{n})^{(1)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\left\|(f_{n})^{(3)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\|g_{n}\|_{B^{s}_{p,r}}=0,$
 but the distance of second component of the difference between the solution sequences $\mathbf{S}_{t}(f_{n}+g_{n})$ and $\mathbf{S}_{t}(f_{n})$
@@
 We should mention that Theorem 1.2 is stronger than that for the Euler equations in [19].
-On the one hand, the non-uniform continuous is established in the more framework of Besov spaces.
+On the one hand, the non-uniform continuous is established in the more general framework of Besov spaces.
 On the other hand, Theorem 1.2 is focused on the second component of data-to-solution map which leads to the non-uniform continuous.
@@
 Theorem 1.3.
-Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3).
-For any $\alpha\in(0,1)$, there exits $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfying
+Assume that $(s,p,r)$ satisfying (1.3).
+For any $\alpha\in(0,1)$, there exists $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.1) satisfies
 $\limsup_{t\to 0^{+}}\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}=+\infty.$
@@
 Theorem 1.4.
-Assume that $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ with $(s,p,r=\infty)$ satisfying (1.3).
-There exits $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) satisfies
+Assume that $(s,p,r=\infty)$ satisfying (1.3).
+There exists $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.1) satisfies
 $\limsup_{t\to 0^{+}}\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,\infty}}\geq\varepsilon_{0}>0.$
@@
 Remark 1.5.
-As mentioned above, (1.2) with $\Omega=0$ is the classical Euler equations. Thus Our main Theorems 1.1-1.4 holds true for the classical Euler equations.
+As mentioned above, (1.1) with $\Omega=0$ is the classical Euler equations. Thus our main Theorems 1.1-1.4 hold true for the classical Euler equations.
 1.4 Strategies to the proof
@@
 The main task is to construct initial data such that the second component of the initial velocity is uniformly bounded in $B^{s}_{p,r}$, and in the same time to have the other component of the initial velocity remain small.
-We first choose two sequences of initial values $u^{n}_{0,1}=f_{n}+g_{n}$ and $u^{n}_{0,2}=f_{n}+g_{n}$, which can generate the approximate solutions $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$, respectively.
+We first choose two sequences of initial values $u^{n}_{0,1}=f_{n}+g_{n}$ and $u^{n}_{0,2}=f_{n}$, which can generate the approximate solutions $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$, respectively.
 Next, we perform perturbation analysis (See Proposition 3.1) to show the second component of the difference between the data-to-solution maps $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$ is bounded below by a positive constant at any later time, which means the solution map is not uniformly continuous.
 Proof of Theorem 1.3.
-In order to prove the failure of Hölder regularity of solutions for the (1.2), we would like to first find initial condition to generate the corresponding solutions to (1.2) satisfying that some positive constant $C$
-$\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}\geq Cn.$
-Our key idea is to decompose the difference between the data-to-solution maps $\mathbf{S}_{t}(u_{0})$ and $\mathbf{S}_{t}(u_{0})$ as
-$\mathbf{S}_{t}(u_{0})-u_{0}=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})-t\mathbf{v}_{0}(u_{0})\quad\text{and}\quad\mathbf{v}_{0}(u_{0})=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0}).$
-There remains two main difficulties that we need to overcome.
+In order to prove the failure of Hölder regularity of solutions for the (1.1), we would like to first find initial condition to generate the corresponding solutions to (1.1) satisfying that
+$\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}\geq C(t)\to+\infty\quad\text{as}\quad t\to 0^{+}.$
+Our key idea is to decompose the difference between the data-to-solution map $\mathbf{S}_{t}(u_{0})$ and initial data $u_{0}$ as
+$\mathbf{S}_{t}(u_{0})-u_{0}=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})-t\mathbf{v}_{0}(u_{0}),$
+here and in what follows we denote $\mathbf{v}_{0}(u_{0}):=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0})$.
+There remains two main difficulties that we need to overcome
   • First, we perform perturbation analysis to prove that $\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})$ is small;
@@
 We construct a new initial data which is somewhat different from Theorem 1.2.
-Finally, by utilizing the commutator estimate and some basic analysis, we hope to obtain the the loss of Hölder regularity of solutions to (1.2).
+Finally, thanks to commutator estimates and some basic analysis, we obtain the loss of Hölder regularity of solutions to (1.1).
 Proof of Theorem 1.4. The main point is to modify the initial data constructed in Theorem 1.3. Following the same procedure as that in Theorem 1.3, we can prove Theorem 1.4.
@@
   • In Section 2, we list some notations and known results which will be used in the sequel.
-  • In Section 3, we show the non-uniform continuous of the solution map (Theorem 1.2).
+  • In Section 3, we show the non-uniform continuity of the solution map (Theorem 1.2).
   • In Section 4, we prove the failure of Hölder regularity of the solution map (Theorem 1.3).
-  • In Section 5, we establish the discontinuous at zero time of the solution map (Theorem 1.4).
+  • In Section 5, we establish the discontinuity at zero time of the solution map (Theorem 1.4).
 2 Preliminaries
@@
   • The inverse Fourier transform allows us to recover $u$ from $\widehat{u}$:
-    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{3}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
-  • We denote the projection
+    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-3}\int_{\mathbb{R}^{3}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
+  • We denote the Leray projection
     $\mathcal{P}:L^{p}(\mathbb{R}^{3})\rightarrow L_{\sigma}^{p}(\mathbb{R}^{3})\equiv\overline{\left\{f\in\mathcal{C}^{\infty}_{0}(\mathbb{R}^{d});{\rm{div}}f=0\right\}}^{\|\cdot\|_{L^{p}(\mathbb{R}^{3})}},\quad p\in(1,\infty),$
     $\mathcal{Q}=\mathrm{Id}-\mathcal{P}.$
-    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant 3}$, where $\mathcal{P}_{ij}\equiv\delta_{ij}+R_{i}R_{j}$ with $\delta_{ij}$ being the Kronecker delta ($\delta_{ij}=0$ for $i\neq j$ and $\delta_{ii}=0$) and $R_{i}$ being the Riesz transform with symbol $-\mathrm{i}\xi_{1}/|\xi|$.
+    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant 3}$, where $\mathcal{P}_{ij}\equiv\delta_{ij}+R_{i}R_{j}$ with $\delta_{ij}$ being the Kronecker delta ($\delta_{ij}=0$ for $i\neq j$ and $\delta_{ii}=0$) and $R_{i}$ being the Riesz transform with symbol $\mathrm{i}\xi_{i}/|\xi|$.
 Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
@@
   • ${\rm{supp}}\;\varphi\subset\left\{\xi\in\mathbb{R}^{3}:3/4\leq|\xi|\leq 8/3\right\}$;
-  • $\varphi(\xi)\equiv 1$ for $24/3\leq|\xi|\leq 3/$ 2;
+  • $\varphi(\xi)\equiv 1$ for $4/3\leq|\xi|\leq 3/2$;
   • $\vartheta(\xi)+\sum_{j\geq 0}\varphi(2^{-j}\xi)=1$ for any $\xi\in\mathbb{R}^{3}$;
@@
 and
-$\forall\,u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{3}),\quad\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
+$\forall\,u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{3}),\quad\dot{\Delta}_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
 where the pseudo-differential operator is defined by $\sigma(D):u\to\mathcal{F}^{-1}(\sigma\mathcal{F}u)$ and $\mathcal{S}^{\prime}_{h}$ is given by
@@
 We recall the definition of the Besov spaces and norms.
-Definition 2.1 ([3]).
-Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the nonhomogeneous and homogeneous Besov spaces, respectively,
-$B^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}:\;\|f\|_{B^{s}_{p,r}}:=\left\|2^{js}\|\Delta_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\geq-1)}<\infty\right\}$
-and
+Definition 2.1 (Nonhomogeneous Besov spaces, see [3]).
+Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the nonhomogeneous Besov spaces
+$B^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}:\;\|f\|_{B^{s}_{p,r}}:=\left\|2^{js}\|\Delta_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\geq-1)}<\infty\right\}.$
+Definition 2.2 (Homogeneous Besov spaces, see [3]).
+Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the homogeneous Besov spaces
 $\dot{B}^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}_{h}:\;\|f\|_{\dot{B}^{s}_{p,r}}:=\left\|2^{js}\|\dot{\Delta}_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\in\mathbb{Z})}<\infty\right\}.$
@@
 Lemma 2.3 ([18]).
-Assume that $(s,p,r)$ satisfies (1.3). Then there exists a constant $C$, depending only on $d,p,r,s$, such that for all $u,v\in B^{s}_{p,r}$ with $\mathrm{div\,}u=\mathrm{div\,}v=0$,
-$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\|u\|_{B^{s}_{p,r}}\|v\|_{B^{s}_{p,r}}$
-and
+Assume that $(s,p,r)$ satisfies (1.3). Then there exists a constant $C$, depending only on $d,p,r,s$, such that for all $u,v\in B^{s}_{p,r}$ with $\mathrm{div\,}u=\mathrm{div\,}v=0$
+$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\left(\|u\|_{C^{0,1}}\|v\|_{B^{s}_{p,r}}+\|v\|_{C^{0,1}}\|u\|_{B^{s}_{p,r}}\right)$
+with $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$. Furthermore, there holds
+$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\|u\|_{B^{s}_{p,r}}\|v\|_{B^{s}_{p,r}},$
 $\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s-1}_{p,r}}\leq C\min\left\{\|u\|_{B^{s-1}_{p,r}}\|v\|_{B^{s}_{p,r}},\,\|v\|_{B^{s-1}_{p,r}}\|u\|_{B^{s}_{p,r}}\right\}.$
@@
 ${\widehat{\phi}(\xi)=}\begin{cases}1,&\text{if}\quad|\xi|\leq\frac{1}{64},\\ 0,&\text{if}\quad|\xi|\geq\frac{1}{8}.\end{cases}$
-We assume that $\widehat{\phi}$ is even and real-valued such that $\theta$ is a real-valued function. Then we can define the new real-valued function
+The inverse Fourier transform allows us to recover $\phi$ from $\widehat{\phi}$, namely, $\phi(x)=\mathcal{F}^{-1}(\widehat{\phi}(\xi))$. Then we can define the new real-valued function
 $\theta(x)=\phi(x_{1})\phi(x_{2})\phi(x_{3}).$
 It is easy to check that $\theta(0)=\phi^{3}(0)>0$ and
-$\mathrm{supp}\ \widehat{\theta}\subset\left\{\xi=(\xi_{1},\xi_{2},\xi_{3})\in\mathbb{R}^{3}:\ |\xi|\leq\frac{1}{2}\right\}.$ (3.6)
+[…]
 Thus for $\lambda\gg 1$ and $T\in\{\sin,\,\cos\}$
-$\mathrm{supp}\ \mathcal{F}\Big{(}\theta(x)T(\lambda x_{1})\Big{)}\subset\left\{\xi\in\mathbb{R}^{3}:\ \lambda-\frac{1}{2}\leq|\xi|\leq\lambda+\frac{1}{2},\ |\xi_{2}|\leq\frac{1}{2},\ |\xi_{3}|\leq\frac{1}{2}\right\}.$ (3.7)
+$\mathrm{supp}\ \mathcal{F}\Big(\theta(x)T(\lambda x_{1})\Big)\subset\left\{\xi\in\mathbb{R}^{3}:\ \lambda-\frac{1}{8}\leq|\xi_{1}|\leq\lambda+\frac{1}{8},\ |\xi_{2}|\leq\frac{1}{8},\ |\xi_{3}|\leq\frac{1}{8}\right\}.$ (3.7)
 For any $p\in[1,\infty]$, there exist two positive constants $C_{1}$ and $C_{2}$ such that
@@
 $g_{n}(x)=2^{-n}\left(-\partial_{2},\partial_{1},0\right)\theta(x).$
+Remark 3.1.
+The choice of $\frac{17}{12}\in(\frac{4}{3},\frac{3}{2})$ is crucial but not unique. In fact, due to (3.7), which implies that
+$\mathrm{supp}\ \mathcal{F}\left(f_{n}\right)\subset\left\{\xi\in\mathbb{R}^{3}:\ \frac{17}{12}2^{n}-1\leq|\xi|\leq\frac{17}{12}2^{n}+1\right\},$
+and notice that $\varphi(2^{-j}\xi)\equiv 1$ in $\left\{\xi\in\mathbb{R}^{3}:\ \frac{4}{3}2^{j}\leq|\xi|\leq\frac{3}{2}2^{j}\right\},$ we have $\mathcal{F}(\dot{\Delta}_{j}f_{n})=\varphi(2^{-j}\cdot)\widehat{f_{n}}=0$ for $j\neq n,$ and thus $\dot{\Delta}_{j}f_{n}=f_{n}$ if $j=n$.
 Lemma 3.1.
@@
 Proof.
-From (3.7) and Bernstein’s inequality (see Lemma 2.1), it follows that the desired results of Lemma 3.1. ∎
+From Remark 3.1 and Bernstein’s inequality (see Lemma 2.1), we obtain the desired results of Lemma 3.1. ∎
 Lemma 3.2.
@@
 $\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,\infty}}$ $=2^{ns}\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{L^{p}}.$ (3.9)
-The definitions of $f_{n}$ and $g_{n}$ tell directly us that
-$2^{ns}g_{n}\cdot\nabla f_{n}^{(2)}$ $=\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
-$\quad+2^{-n}\phi(x_{1})\phi^{\prime}(x_{1})\sin\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
-$\quad-2^{-2n}\phi(x_{1})\phi^{\prime\prime}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
-$\quad+2^{-2n}\left[\phi^{\prime}(x_{1})\right]^{2}\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3}).$
-Then, we have
+The definitions of $f_{n}$ and $g_{n}$ directly tell us that
+$2^{ns}g_{n}\cdot\nabla f_{n}^{(2)}$ $=-\left(\frac{17}{12}\right)^{2}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})+\mathbf{R}.$
+For the remainder terms $\mathbf{R}$, noticing that at most one partial derivative falls on cosine function, then $\|\mathbf{R}\|_{L^{p}}\leq C2^{-n}$. Thus, we have
 $2^{ns}\|g_{n}\cdot\nabla f_{n}^{(2)}\|_{L^{p}}\geq\left\|\phi^{2}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(\mathbb{R})}\|\phi\|^{2}_{L^{2p}(\mathbb{R})}\left\|\phi(x)\phi^{\prime}(x)\right\|_{L^{p}(\mathbb{R})}-C2^{-n},$
@@
 $\|\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}\leq Ct^{2}\mathbf{E}(u_{0}),$ (3.11)
-here and in what follows we denote $\mathbf{v}_{0}(u_{0}):=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0})$ and
+here and in what follows we denote
 $\mathbf{E}(u_{0})$ $:=1+\|u_{0}\|_{B^{s-1}_{p,r}}\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$
@@
 $\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
-$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s-1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s-1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
-$\leq C\int^{t}_{0}\left(\|u\|_{B^{s}_{p,r}}\|u\|_{B^{s-1}_{p,r}}+\|u\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
-$\leq Ct\|u_{0}\|_{B^{s-1}_{p,r}},$ (3.13)
+$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s-1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s-1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
+$\leq$ $~C\int^{t}_{0}\left(\|u\|_{B^{s}_{p,r}}\|u\|_{B^{s-1}_{p,r}}+\|u\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
+$\leq$ $~Ct\|u_{0}\|_{B^{s-1}_{p,r}},$ (3.13)
 $\|u(t)-u_{0}\|_{B^{s}_{p,r}}\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s}_{p,r}}\mathrm{d}\tau$
-$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s}_{p,r}}\right)\mathrm{d}\tau$
-$\leq Ct\left(1+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+1}_{p,r}}\right),$ (3.14)
+$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s}_{p,r}}\right)\mathrm{d}\tau$
+$\leq$ $~Ct\left(1+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+1}_{p,r}}\right),$ (3.14)
 $\|u(t)-u_{0}\|_{B^{s+1}_{p,r}}\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
-$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s+1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s+1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s+1}_{p,r}}\right)\mathrm{d}\tau$
-$\leq Ct\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$ (3.15)
+$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s+1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s+1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s+1}_{p,r}}\right)\mathrm{d}\tau$
+$\leq$ $~Ct\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$ (3.15)
 Using the Mean Value Theorem once again, we obtain that
 $\|u(t)-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}\leq\int^{t}_{0}\|\partial_{\tau}u+\mathbf{v}_{0}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau$
-$\leq\int^{t}_{0}\|\mathcal{Q}(u)-\mathcal{Q}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\cdot\nabla u-u_{0}\cdot\nabla u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau$
-$\quad+\Omega\int^{t}_{0}\|\mathcal{P}\left(e_{3}\times(u-u_{0})\right)\|_{B^{s}_{p,r}}\mathrm{d}\tau$
-$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\|u(\tau)\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
-$\quad\ +\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\|u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
-$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\|u_{0}\|_{B^{s+1}_{p,r}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
-$\quad\ +\|u_{0}\|_{L^{\infty}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\mathrm{d}\tau.$ (3.16)
+$\leq$ $~\int^{t}_{0}\|\mathcal{Q}(u)-\mathcal{Q}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\cdot\nabla u-u_{0}\cdot\nabla u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau$
+$+\Omega\int^{t}_{0}\|\mathcal{P}\left(e_{3}\times(u-u_{0})\right)\|_{B^{s}_{p,r}}\mathrm{d}\tau$
+$\lesssim$ $~\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\|u(\tau)\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
+$+\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\|u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
+$\lesssim$ $~\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\|u_{0}\|_{B^{s+1}_{p,r}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
+$+\|u_{0}\|_{L^{\infty}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\mathrm{d}\tau.$ (3.16)
 Plugging (3)-(3) into (3) and using the fact that $B^{s-1}_{p,r}(\mathbb{R}^{3})\hookrightarrow L^{\infty}(\mathbb{R}^{3})$, yields the desired result of Proposition 3.1. ∎
@@
 We decompose the solution maps $\mathbf{S}_{t}(u^{n}_{0})$ and $\mathbf{S}_{t}(f_{n})$ as follows
-$\mathbf{S}_{t}(u^{n}_{0})$ $=\mathbf{S}_{t}(u^{n}_{0})-u^{n}_{0}+t\mathbf{v}_{0}(u^{n}_{0})-(f_{n}+g_{n})-t(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})$
+$\mathbf{S}_{t}(u^{n}_{0})$ $=\mathbf{S}_{t}(u^{n}_{0})-u^{n}_{0}+t\mathbf{v}_{0}(u^{n}_{0})+(f_{n}+g_{n})-t(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})$
 $\quad+t\mathcal{Q}\big[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})\big]-t\mathcal{P}(\Omega e_{3}\times(f_{n}+g_{n})),$
-$\mathbf{S}_{t}(f_{n})$ $=\mathbf{S}_{t}(f_{n})-f_{n}+t\mathbf{v}_{0}(f_{n})-f_{n}-tf_{n}\cdot\nabla f_{n}$
+$\mathbf{S}_{t}(f_{n})$ $=\mathbf{S}_{t}(f_{n})-f_{n}+t\mathbf{v}_{0}(f_{n})+f_{n}-tf_{n}\cdot\nabla f_{n}$
 $\quad+t\mathcal{Q}(f_{n}\cdot\nabla f_{n})-t\mathcal{P}(\Omega e_{3}\times f_{n}).$
 Then, by Proposition 3.1, we deduce that
-$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$ $\geq t\left\|\left[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\right]^{(2)}\right\|_{B^{s}_{p,r}}-t\|\mathcal{P}(\Omega e_{3}\times g_{n})\|_{B^{s}_{p,r}}$
-$\quad-t\left\|\mathcal{Q}\big{[}(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\big{]}\right\|_{B^{s}_{p,r}}-\|g_{n}\|_{B^{s}_{p,r}}-Ct^{2}$
-$\geq t\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,r}}-Ct^{2}-C2^{-n},$ (3.17)
+$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$
+$\geq$ $~t\left\|\left[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\right]^{(2)}\right\|_{B^{s}_{p,r}}-t\|\mathcal{P}(\Omega e_{3}\times g_{n})\|_{B^{s}_{p,r}}$
+$-t\left\|\mathcal{Q}\big[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\big]\right\|_{B^{s}_{p,r}}-\|g_{n}\|_{B^{s}_{p,r}}-Ct^{2}$
+$\geq$ $~t\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,r}}-Ct^{2}-C2^{-n},$ (3.17)
 where we have used the facts from Lemma 2.3 and Lemma 3.1
@@
 Define the divergence-free vector field $u_{0}$ as above.
-Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some positive constant $c_{0}$ such that
+Then there exists some sufficiently large $n\in\mathbb{N}$ and some positive constant $c_{0}$ such that
 $\|u_{0}\|_{B^{s}_{p,r}}\thickapprox\|u_{0}\|_{\dot{B}^{s}_{p,r}}\thickapprox 1,$ (4.18)
@@
 Proof.
-Due to the simple fact (3.7), it holds that for some sufficiently large $n\in\mathbb{Z}^{+}$
+Due to the simple fact (3.7), it holds that for some sufficiently large $n\in\mathbb{N}$
 $\Delta_{n}u_{0}^{(1)}=-\partial_{2}a_{n}\quad\text{and}\quad\Delta_{n}u_{0}^{(2)}=\partial_{1}a_{n},$
@@
 $-n^{2}2^{n(s+1)}u_{0}^{(1)}{\partial_{1}\partial_{1}}a_{n}$ $=\left(\frac{17}{12}2^{n}\right)^{2}\sum_{k=3}^{\infty}k^{-2}2^{-k(s+1)}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{k}x_{1}\right)$
-×$\quad\times\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi^{2}(x_{3})\phi(x_{2})\phi^{\prime}(x_{2})+\text{Remainder terms}.$
-For the Remainder terms, noticing that at most one partial derivative falls on cosine function, thus we find easily that
+$\quad\times\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi^{2}(x_{3})\phi(x_{2})\phi^{\prime}(x_{2})+\text{Remainder terms}.$
+For the Remainder terms, noticing that at most one partial derivative falls on cosine function, thus we easily find that
 $\left\|\text{Remainder terms}\right\|_{L^{p}(\mathbb{R}^{3})}\leq C2^{n}.$
@@
 $\left|\sum_{k=3}^{\infty}k^{-2}2^{-k(s+1)}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{k}x_{1}\right)\right|\geq\frac{\phi^{2}(0)}{2}\sum\limits^{\infty}_{k=3}k^{-2}2^{-k(s+1)}=:c_{0}>0.$
-Thus we have for some sufficiently large $n\in\mathbb{Z}^{+}$
-$n^{2}2^{n(s+1)}\left\|u_{0}^{(1)}{\bf\partial_{1}\partial_{1}}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\geq c_{0}2^{2n}-C2^{n}.$ (4.21)
+Thus we have for some sufficiently large $n\in\mathbb{N}$
+$n^{2}2^{n(s+1)}\left\|u_{0}^{(1)}\partial_{1}\partial_{1}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\geq c_{0}2^{2n}-C2^{n}.$ (4.21)
 By easy computations, we have
 $n^{2}2^{n(s+1)}\left\|u_{0}^{(2)}\partial_{1}\partial_{2}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\leq C2^{n}.$ (4.22)
-Noticing that (4.20), from (4.21) and (4.22), we obtain that for some sufficiently large $n\in\mathbb{Z}^{+}$
+Noticing that (4.20), from (4.21) and (4.22), we obtain that for some sufficiently large $n\in\mathbb{N}$
 $n^{2}2^{n(s+1)}\left\|u_{0}\cdot\nabla\Delta_{k}u_{0}^{(2)}\right\|_{L^{p}(\mathbb{R}^{3})}\geq c_{0}2^{2n}-C2^{n},$
@@
 $\|\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})\|_{{B}^{s-2}_{p,r}}\leq\int^{t}_{0}\|\partial_{\tau}u+\mathbf{v}_{0}\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
-$\lesssim\int^{t}_{0}\|\mathcal{P}\left(u\cdot\nabla u-u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times(u-u_{0})\right)\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
-$\lesssim\int^{t}_{0}\left(\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
-$\lesssim\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$
+$\lesssim$ $~\int^{t}_{0}\|\mathcal{P}\left(u\cdot\nabla u-u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times(u-u_{0})\right)\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
+$\lesssim$ $~\int^{t}_{0}\left(\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
+$\lesssim$ $~\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$
 We complete the proof of Proposition 4.1. ∎
@@
 $\approx 2^{{n(s-1)}}\|\Delta_{n}(\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i})\|_{L^{p}}$
-$\lesssim\|\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}\|_{B^{s-1}_{p,\infty}}$
-$\lesssim\|u_{0}\|^{2}_{B^{s}_{p,\infty}}\lesssim 1.$
+$\lesssim\|\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}\|_{B^{s-1}_{p,\infty}}\lesssim\|u_{0}\|^{2}_{B^{s}_{p,\infty}}\lesssim 1.$
 Then, using Lemma 4.1, we obtain from (4) that
@@
 $u_{0}(x):=\left(-\partial_{2},\partial_{1},0\right)\sum\limits^{\infty}_{n=3}b_{n}(x),\quad\text{where}$
-$b_{n}(x):=2^{-ns}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
-Following similar procedure as that in Theorem 1.3. Then, we have
+$b_{n}(x):=2^{-n(s+1)}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
+Following similar procedure as that in Theorem 1.3, then we have
 $\left\|\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{B^{s}_{p,\infty}}$ $\geq 2^{{ns}}\left\|\Delta_{n}\left(\mathbf{S}_{t}(u_{0})-u_{0}\right)^{(2)}\right\|_{L^{p}}=2^{{ns}}\left\|\Delta_{n}\big(t\mathbf{v}_{0}(u_{0})-\mathbf{w}\big)^{(2)}\right\|_{L^{p}}$
@@
 Math.
-Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces.
+Phys., 296, 2010] to the more general framework of Besov spaces; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces.
 In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero.
 To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force.
-In this paper, our contributions on the continuous properties for the 3D incompressible rotating Euler equationsare lies in three aspects: by constructing some proper, new and different initial conditions, we showed the non-uniform continuous of the solution map (Theorem 1.2), proved the failure of Hölder regularity of the solution map (Theorem 1.3) and established the discontinuous at $t=0$ of the solution map (Theorem 1.4).
+In this paper, our contribution on the continuous properties for the 3D incompressible rotating Euler equations, unfold in three aspects: by constructing some proper, new and different initial conditions, we showed the non-uniform continuous of the solution map (Theorem 1.2), proved the failure of Hölder regularity of the solution map (Theorem 1.3) and established the discontinuous at $t=0$ of the solution map (Theorem 1.4).
 Future directions.
@@
 We will investigate this interesting problem in the future work.
-Acknowledgments
+Declarations
+Availability of data and materials
+No data was used for the research described in the article.
+Conflict of interest
+The authors declare that they have no conflict of interest.
+Funding
+The authors would like to thank the anonymous referees for valuable comments and suggestions which greatly improved the presentation of this paper.
 J.
-Li is partially supported by the National Natural Science Foundation of China (No.
+Li is supported by the National Natural Science Foundation of China (No.
 12161004), Innovative High end Talent Project in Ganpo Talent Program (No.
 gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No.
-20232BCJ23009).
-Y.
-Yu is partially supported by the National Natural Science Foundation of China (No.
-12101011).
+20232BCJ23009), Jiangxi Provincial Natural Science Foundation (No.
+20252BAC210004).
 N.
@@
 20243BCE51079).
-Declarations
-Data Availability No data was used for the research described in the article.
-Conflict of interest The authors declare that they have no conflict of interest.